OpenStudy (anonymous):

You have 20g of cobalt. How many atoms of cobalt is this?

OpenStudy (xishem):

This is similar to the other question you asked. First convert from grams of Co to mols of Co, and then from mols of Co to atoms of Co using the knowledge that there are 6.022*10^23 of something in 1mol of something.

OpenStudy (anonymous):

I got 5.65x10^21 atoms

OpenStudy (xishem):

Well, let's see... \[20g(Co)*\frac{1mol(Co)}{58.933g(Co)}=0.33937mol(Co)\]\[0.33937mol(Co)*\frac{6.022*10^{23}atoms(Co)}{1mol(Co)}=2.0437atoms(Co)=2.0*10^{23}atoms(Co)\]Do you see where you made your mistake?

OpenStudy (xishem):

In the second set of equations, it should be...\[2.0437*10^{23}atoms(Co)=2.0*10^{23}atoms(Co)\]

OpenStudy (xishem):

Just remember to always make sure that your units make sense. If you follow your units, 90% of the time (especially in entry-level chemistry), you will be lead in the correct direction.